OpenStudy (anonymous):

8 7 ---- = --- Solve for x x-7 2

sam (.sam.):

(8)/(x-7)=(7)/(2) 8*2=7*(x-7) 7*(x-7)=8*2 7(x-7)=8*2 7(x-7)=16 7x-49=16 7x=65 x=(65)/(7)